1) A 3300-pF capacitor is connected in series to a 43.0-μH coil of resistance 7.30 Ω .

A)What is the resonant frequency of this circuit?

2) The variable capacitor in the tuner of an AM radio has a capacitance of 2200 pF when the radio is tuned to a station at 560 kHz .

Part A

What must be the capacitance for a station at 1700 kHz ?

Part B

What is the inductance (assumed constant)?

3) An LRC circuit has L = 14.8 mH and R = 3.10 Ω .

Part A

What value must C have to produce resonance at 4600 Hz ?

Part B

What will be the maximum current at resonance if the peak external voltage is 140 V ?
Answer #1

ACCORDING TO THE CHEGG GUIDELINES I AM ALLOWED TO ANSWER ONLY (FIRST ) ONE QUESTION AT A TIME.

1) f = 1/2* 3.14 * sqrt ( LC ) = 1/2 * 3.14 * SQRT ( 3300 * 10 ^ -12 * 43.0 * 10 ^ -6 )

F = 422716.4596 HZ OR 4.22716 * 10 ^ 5 HZ =================================ANSWER)